OpenStudy (rosales0228):

GIVING MEDAL!!!!!!! Which statement is true? A polynomial of degree n has at least n – 1 turning points. A polynomial of degree n has at least n – 2 turning points. A polynomial of degree n has at most n – 1 turning points. A polynomial of degree n has at most n – 2 turning points.

OpenStudy (chrisisdope1):

A

OpenStudy (rosales0228):

it was wrong:(

OpenStudy (chrisisdope1):

Or C

OpenStudy (acxbox22):

it is C

OpenStudy (acxbox22):

a linear equation has a degree of 1 n-1=1-1=0 0 turning points is true because a line doesnt have a turning point

OpenStudy (chrisisdope1):

Sorry I Got Them Mixed Up Or Summin